# On Your Own: Pro-level Autonomous Drone Racing in Uninstrumented Arenas

### v1

@@ VIII Discussion
State-of-the-art research had shown that autonomous drone racing can outperform professional pilots in controlled, instrumented arenas [4]. We set ourselves out to answer the question of whether such level of performance could also be achieved in uninstrumented settings—as this will be crucial for the real-world deployment of these methods.
@@
In the instrumented arena experiments (Fig. 6), we observed different critical maneuvers for the human pilots and the autonomous drone. Star23467 and Ion FPV reported the Split-S (Fig. 7) maneuver as problematic for human pilots due to the constrained physical space, as the Split-S was placed at the end of the arena, near a wall. All pilots consistently identified the spiral section of the track as a pivotal segment for performance. For the autonomous system, the hairpin turn between gates 2 and 3 was a major limiting factor as it became a failure point when trying to execute trajectories generated with a less conservative TWR.
@@
As a further demonstration of skill transferability, our autonomous system was also deployed in two additional uninstrumented venues—featuring novel race tracks—with only minimal tuning and limited on-site testing. Specifically, we only used 80 additional images for fine-tuning (Fig. 4), we made adjustments to the camera exposure and gain (Sec. IV-B), and we generated new time-optimal trajectories (Sec. IV-C1). These deployments were showcased in public demonstrations at IROS 2024 and the 2024 Abu Dhabi F1 Grand Prix, where the autonomous drone outperformed the professional pilot Ion FPV in both time-trial and head-to-head races. Notably, the F1 track posed the additional challenge of outdoor lighting conditions. Videos of both demonstrations are available at: youtube.com/watch?v=SNw-zXgv_vA.
@@ IX Conclusions
In this letter, we presented a drone racing autonomy stack—comprising of vision, state estimation, and control modules—that is capable of pro-pilot performance in uninstrumented arenas, i.e., without access to ground-truth data to fine-tune the state estimation and control modules. In particular, we showed that our drone outperformed professional pilots in an instrumented setting and was competitive against a champion pilot in the uninstrumented one. In our future work, we will replace the stereo-camera with a monocular one (bringing our stack closer to its human counterpart).
X Acknowledgments
We thank the pilots, Thomas Kund (Star23467), Krutharth M. C. (Ion FPV), and Minchan Kim (MCK), for their insight and invaluable suggestions.

### v2

@@ VIII Discussion
State-of-the-art research has shown that autonomous drone racing can outperform professional pilots in controlled, instrumented arenas [4]. We set ourselves out to answer the question of whether such a level of performance could also be achieved in uninstrumented settings—as this will be crucial for the real-world deployment of these methods.
@@
The design choices in Sec. IV and their rigorous integration were critical to achieve the demonstrated champion-level performance. Notably, the smooth state estimate produced by the dual-stage filtering (Sec. IV-B) enabled the aggressive MPC tuning required for lap-time optimization; until the introduction of such filtering, the resulting state discontinuities had led to instability and crashes. Furthermore, the state predictor used for delay compensation (Sec. IV-C2) unlocked 3D maneuvers at more sustained speed. Together, these improvements helped us cross the boundary between achieving pro-level performance and champion-level performance, corresponding to an advantage of approximately 2 seconds on the reference tracks.
In the instrumented arena experiments (Fig. 6), we observed different critical maneuvers for the human pilots and the autonomous drone. Star23467 and Ion FPV reported the Split-S (Fig. 7) maneuver as problematic for human pilots due to the constrained physical space, as the Split-S was placed at the end of the arena, near a wall. All pilots consistently identified the spiral section of the track as a pivotal segment for performance. For the autonomous system, the hairpin turn between gates 2 and 3 was a major limiting factor, as it became a failure point when trying to execute trajectories generated with a less conservative TWR.
@@
As a further demonstration of skill transferability, our autonomous system was also deployed in two additional uninstrumented venues—featuring novel race tracks—with only minimal tuning and limited on-site testing. Specifically, we only used 80 additional images for fine-tuning (Fig. 4), we made adjustments to the camera exposure and gain (Sec. IV-B), and we generated new time-optimal trajectories (Sec. IV-C1). These deployments were showcased in public demonstrations at IROS 2024 and the 2024 Abu Dhabi F1 Grand Prix, where the autonomous drone outperformed the professional pilot Ion FPV in both time-trial and head-to-head races. Notably, the F1 track posed the additional challenge of outdoor lighting conditions. Videos of both demonstrations are included in the multimedia material.
@@ IX Conclusions
In this letter, we presented a drone racing autonomy stack—comprising vision, state estimation, and control modules—that is capable of pro-pilot performance in uninstrumented arenas, i.e., without access to ground-truth data to fine-tune the state estimation and control modules. In particular, we showed that our drone outperformed professional pilots in an instrumented setting and was competitive against a champion pilot in the uninstrumented one. In our future work, we will replace the stereo-camera with a monocular one (bringing our stack closer to its human counterpart).
Acknowledgments—We thank the pilots, Thomas Kund (Star23467), Krutharth M. C. (Ion FPV), and Minchan Kim (MCK), for their insight and invaluable suggestions.